The East Hartford WWTP discharges 150 MGD of treated wastewater into the CT river. The treated was tested in the lab and found to have aBOD5 of 17.5 mg/L and an ultimate BOD (Lo) of 92 mg/L at 20oC....

  1. The East Hartford WWTP discharges 150 MGD of treated wastewater into the CT river. The treated was tested in the lab and found to have aBOD5 of 17.5 mg/L and an ultimate BOD (Lo) of 92 mg/L at 20oC.
  1. Find the reaction rate K, for the treated wastewater.
  2. The CT river has a flow of 250 m2/s and an ultimate BOD (Lo) of 6 mg/L. the DO of the river is 4.5 mg/L and the DO of the wastewater is0.2 mg/L. find the DO and the initial ultimate BOD (Lo) of the downstream mixture if the temperature of the water is 5oC.

  1. Find the initial deficit (Do) in the DO after mixing. Saturated dissolved oxygen values can be found in the textbook appendix.

  1. The deoxygenation rate in the river (after mixing) is kd =0.3d-1 . the reaction rate in the river is ks = 0.5 d-1. If the river is traveling at 0.25 m/s, what will the deficit, D, be 4.32 km from the point where the treated wastewater enters the river?
Answer #1

Answers:

a) BOD5 = UBOD * ( 1 - e^(-k*t))

17.5 = 92*( 1 - e^(-k*5)

k = 0.042 day^-1

b) UBOD = (Q1*UBOD1 + Q2*UBOD2) / (Q1 + Q2)

1 MGD = 0.043813 m3/s

150 MGD = 0.043813 * 150 m3/s

UBOD =( 0.043813 * 150 * 92 + 250*6 ) / (0.043813 * 150 + 250)

= 8.2 mg/L

DO = ( 0.043813 * 150 * 0.2 + 250*4.5 ) / (0.043813 * 150 + 250)

= 4.39 mg/L

deficit = DOsat - DO

critical travel time = Tc = 4.32 * 1000 / 0.25

= 17280 s

= 0.2

critical deficit = Dt =( kd * UBOD / (ks - kd)) * [ e^(-kd*t) - e^(ks*t) ] + [ Da*e^-ks*t]

assuming initial deficit Da = 0

Dt = ( 0.3 * 8.2 / (0.5 - 0.3)) * [ e^(-0.3*0.2) - e^(0.5*0.2) ]

=- 2 mg/L + Da